# Towards an Understanding of Context Utilization in Code Intelligence

### v1

@@ 3.3. Inclusion and Exclusion Criteria
After identifying studies that align with our research objectives, it is necessary to filter out unqualified studies. To achieve this, we apply inclusion and exclusion criteria to assess the quality of candidate studies. This evaluation ensures that each retained research primarily focuses on using context to address one type of code intelligence task. The inclusion and exclusion criteria are as follows.
✓The paper must be published between September 1, 2007 and August 31, 2024.
@@ 3.4. Data Extraction and Collection
After filtering out irrelevant and duplicate papers, as shown in Figure 1, we extract relevant data and analyze it to address our four research questions. For RQ1, we record the year, type, publication venue, primary contributions, and task focus of each study. These details are vital for mapping the evolution of the research landscape, revealing historical trends, and understanding how various publication venues have influenced the development of context-aware code intelligence. For RQ2, we examine the types of context used and analyze trends in their application across seven code intelligence tasks. Collecting these data points is instrumental in identifying the diversity of context usage and in assessing its potential application scenarios. For RQ3, we review the context preprocessing methods and the context-aware models developed in papers introducing new techniques. This data collection sheds light on the methodological innovations and technical challenges involved in effectively leveraging context, which is crucial for advancing model development. For RQ4, we collate relevant metrics and available context types in datasets, evaluating open-source conditions in context-relevant papers. These aspects are essential for future evaluation strategies and research reproducibility. To ensure data accuracy, two researchers independently double-check the results.
@@ 4. RQ1: What Are the Trends and Distributions of the Publications Studied?
To understand the development trends of papers that utilize context in code intelligence, we analyze the trends and distributions of the reviewed publications. Our analysis focuses on three key aspects: the publication dates, the venues where the studies were published, and the main contributions of the primary studies.
@@ 4.1. Trends of Publication Year
[…]
[Short description]Number of Papers Published Per Year
@@
Although the concept of “context” had been explored in natural language processing tasks before 2007 (Porzel and Gurevych, 2003), to the best of our knowledge, we did not come across any papers prior to 2007 that used context to address code intelligence tasks during our investigation. Therefore, our analysis focuses on relevant papers published from 2007 to 2024. Figure  2 illustrates the publication trend of works employing context over the past eighteen years. To provide a clearer view of the trend in context utilization within code intelligence, we divide the research of these papers presented in these papers into three distinct phases.
@@
The second phase is from 2019 to 2022, during which more relevant work emerges. The average number of published papers per year rises to 9.5. This increase is related to earlier work that laid the foundation for context utilization and developing deep learning models, such as the transformer architecture (Vaswani et al., 2017). At this stage, all seven types of code intelligence tasks begin to investigate how to utilize contextual information fully.
The third phase is from 2023 to 2024. In this phase, ChatGPT (Wu et al., 2023) starts a wave of large language models, whose outstanding performance drives code intelligence tasks to further explore the utilization of context based on large language models. During this phase, the number of papers increases substantially, with 31 papers published in 2023 and 49 in 2024 (as of August). Except for the clone detection task, all tasks show substantial growth in the number of papers. We suspect that clone detection may be less dependent on large language models.
[…]
[Short description](Left) Number of Publications Per Type. (Right) Number of Publications Per Venue.
@@ 4.3. Types of Main Contributions of the Studied Papers
[…]
[Short description](Left) Number of Papers Per Type of Main Contribution. (Right) Number of Papers Per Tasks.
@@
We simultaneously analyze the proportion of each type of code intelligence task, which is relatively distributed as Figue  4 shows. Top-3 code intelligence tasks are defect detection (19.9%), code completion (19.2%), and code summarization (16.4%), while code generation occupies 15.8% and program repair contains 13.7%. Commit message generation and clone detection take the minor parts, 8.2%, and 6.8%, respectively.
@@ 5. RQ2: How Are Contexts Categorized in Code Intelligence Tasks?
To categorize context in code intelligence, we divide it into two types based on how it is obtained: direct context and indirect context. Direct context refers to information that can be accessed without further processing, such as source code, API documentation, and commit messages. Indirect context, however, requires additional processing or transformation, examples of which include abstract syntax trees (ASTs), control flow graphs (CFGs), and compiler-generated information. This section will explore the different context categories and examine the trends in their usage within code intelligence tasks.
@@ 5.1. Types of Context
As shown in Figure  5, we summarize all types of contexts into a tree diagram. The number after the name of the context represents how many new technique papers utilize this kind of context, while benchmark, case study, and empirical papers are excluded. The combined count of direct context utilization (97 papers) and indirect context utilization (51 papers) exceeds the total number of papers (122 papers) because 26 papers utilize both direct and indirect contexts. Overall, direct contexts are more prevalent than indirect contexts. However, researchers discovered that transforming direct context into various indirect contexts in specific scenarios can improve model performance and efficiency. For example, Zhang et al. (Zhang et al., 2023b) propose a graph neural network based on control flow graphs to detect condition-related bugs, and this model significantly outperforms the state-of-the-art methods. Li et al. (Li, 2020) embed abstract syntax trees, data flow graphs, and program dependency graphs to capture local and global context for bug detection. These representations relatively improve over the other representations up to 206% in accuracy.
[…]
[Short description]Types of Context Used in the Seven Studied Code Intelligence Tasks
@@ 5.1.1. Direct Context
Eighty papers adapt code-based contexts, which are the most numerous of all forms of contexts, including sixty-eight papers for source code and seventeen papers for code diffs. Source code, the most popular form of context, constructs the basis for program execution. Mining and extracting source code context mitigates hallucination and provides more accurate answers by offering models with specific domain knowledge in repositories (Nashid et al., 2024; Zhang et al., 2023a). Code diff context presents changes in source code over time or across versions, which is suitable for providing more fine-grained information compared with source code context in tasks such as defect detection (Youm et al., 2017) and commit message generation (Li et al., 2024a).
@@
[…]
[Short description]Number of Context Utilized in Each Code Intelligence Task
Figure 6. Number of Context Utilized in Each Code Intelligence Task
@@ 5.1.2. Indirect Context
Indirect context can be categorized into three types: tree-based, graph-based, and sequence-based contexts. Tree-based context includes AST (24 papers), and the construction of AST showcases the structure information of code. The graph-based context contains CFG (9 papers), DFG (6 papers), PDG (6 papers), CPG (5 papers), and UML (2 papers), demonstrating important information in source code such as data flow, control flow, logic, and structure. Sequence-based context includes compilation information (9 papers) and IDE information (2 papers). Compilation information offers insight into the dynamics of code execution, and IDE information provides information such as programmers’ usage patterns.
@@ 5.2. Trends of Context
Besides, we find that source code is the most popular form of context, utilized by all seven code intelligence tasks, followed by AST (utilized by six tasks) and CFG (utilized by five tasks). In all 13 kinds of context, defect detection employs the widest variety of context (9 types), while program repair and clone detection both select eight kinds of context, respectively. However, some kinds of context are less common. Commit message and IDE information are utilized twice by commit message generation (Huang et al., 2017; Eliseeva et al., 2023) and code completion (Li et al., 2024d; Bibaev et al., 2022), respectively. UML is selected by code generation (E. V. and Samuel, 2019) and code summarization (Wang et al., 2021a).
The variation in the amount of context utilized across different code intelligence tasks is due to how well the context information aligns with the specific requirements of each task. For example, obtaining functional details needed for code completion can be particularly challenging from UML diagrams, which are not always well-suited.
@@ 6. RQ3: What Methods Are Used to Preprocess and Model Context?
While context provides valuable information, it must be carefully processed and modeled to be effectively used in code intelligence tasks. Preprocessing refers to the steps taken to clean, structure, or transform raw context data into a usable format, such as tokenizing code or splitting identifiers. Modeling involves the techniques used to integrate this processed context into machine learning or deep learning models, allowing the context to inform and improve task performance. This section examines how researchers process context, focusing on both the preprocessing methods and the modeling strategies employed in different code intelligence tasks.
@@ 6.1. Context Preprocess
To intuitively understand the effect of context preprocessing in each paper, we categorize them into four types of operators: splitting, irrelevance removal, unifying, and other operator. The detail of the preprocessing steps is shown in Table  3.
@@ 6.1.1. Splitting
Byte pair encoding (BPE) (Gage, 1994) is a widely used splitting method in natural language processing. BPE reduces the number of tokens by iteratively converting frequently occurring data into new units, alleviating the training burden while maintaining data accuracy. For example, Liu et al. (Liu et al., 2023a) employ BPE method to segment both natural language and source code data, facilitating the generation of feature embeddings. This technique can be applied in code intelligence tasks to process code-relevant context, such as API documentation (Zan et al., 2023, 2022a; Zhou et al., 2023), compiler information (Agrawal et al., 2024), code comments (Liu et al., 2023a), source code contexts (Ding et al., 2023b; Shrivastava et al., 2023b; Zhang et al., 2023a).
Programming languages are highly structured and adhere to stricter syntax rules than natural language descriptions. To facilitate program comprehension, programmers often adopt CamelCase or SnakeCase as naming conventions for coding identifiers, such as variables, functions, and classes. These conventions offer practical guidelines for splitting identifiers: CamelCase identifiers can be divided by detecting transitions between lowercase and uppercase letters, while SnakeCase identifiers are split at underscores (’_’). Both techniques are widely used and effective in preprocessing source code context (Lutellier et al., 2020; Wang et al., 2021a; Bansal et al., 2023c). For example, Wang et al. (Wang et al., 2021a) employ the CamelCase rule to split class names in UML, while Bansal (Bansal et al., 2023c) utilize the CamelCase rule to split cross-file source code tokens.
Additionally, segmentation rules based on non-alphabetic (Washio and Miyao, 2022; Xu et al., 2022; Gao et al., 2023) or non-numeric symbols (Washio and Miyao, 2022; Shrivastava et al., 2023a) can serve as supplementary approaches for segmenting both code and natural language content. These methods are widely used due to their simplicity, as they do not rely on complex lexicons or language models and serve as an initial approach for token segmentation.
Another approach is to use a tokenizer from an established library or provided by language models. Examples of established libraries include NLTK   [^1] https://github.com/nltk/nltk (Lin et al., 2021; Bangcharoensap et al., 2012), Javalang   [^2] https://github.com/c2nes/javalang (Zhang et al., 2023b) and Pygments   [^3] https://github.com/pygments/pygments (Zügner et al., 2021). Examples of tokenizers provided by language models include those from StarCoder (Phan et al., 2024; Cheng et al., 2024; Liu et al., 2024b), GPT-4 (Xie et al., 2024), and RoBERTa (Wang et al., 2024f). A tokenizer segments continuous text into discrete tokens, assigning each token a unique index, which is then mapped to token embeddings for further processing.
@@
Splitting BPE (Agashe et al., 2019; Agrawal et al., 2024; Zan et al., 2023, 2022a; Tang et al., 2023; Zhou et al., 2023; Ding et al., 2023b; Shrivastava et al., 2023b; Zhang et al., 2023a; Liu et al., 2023a; Ding et al., 2023a; Qiu et al., 2024) 12
Camel_Case (Lutellier et al., 2020; Lin et al., 2022; Saha et al., 2013; Yu et al., 2020; Shahbazi and Fard, 2023; Shahbazi et al., 2021; Guo et al., 2023a; Shrivastava et al., 2023b; Wang et al., 2021a; Bansal et al., 2023c, a; Zügner et al., 2021; Huang et al., 2017; Lin et al., 2021; Bangcharoensap et al., 2012; Youm et al., 2015; Swe and Oo, 2019; Zhou et al., 2012; Jiang et al., 2018; Wang and Lo, 2014; Gao et al., 2023; Wong et al., 2015; Motwani and Brun, 2023; Youm et al., 2017; Zhang et al., 2021) 25
Snake_Case (Lin et al., 2022; Yu et al., 2020; Shahbazi and Fard, 2023; Shahbazi et al., 2021; Guo et al., 2023a; Shrivastava et al., 2023b; Wang et al., 2021a; Bansal et al., 2023c, a; Zügner et al., 2021; Zhang et al., 2021) 11
@@
Base on Tokenizer (Zügner et al., 2021; Lin et al., 2021; Zhang et al., 2023b; Bangcharoensap et al., 2012; Phan et al., 2024; Xie et al., 2024; Wang et al., 2024g; Cheng et al., 2024; Wang et al., 2024f; Liu et al., 2024b; Deng et al., 2024) 11
@@
Punctuations Filtering (Shahbazi and Fard, 2023; Shahbazi et al., 2021; Wang et al., 2021a; Bansal et al., 2023a; Koyuncu et al., 2019; Matmti and Fard, 2023; Wang and Lo, 2014; Motwani and Brun, 2023; Youm et al., 2017; Zhang et al., 2021) 10
Comment Removal (Zügner et al., 2021; Guo et al., 2024a; Imani et al., 2024) 1
@@
Symbol Removal (Guo et al., 2024a) 1
Unifying Lowercase (Shahbazi and Fard, 2023; Shahbazi et al., 2021; Wang et al., 2021a; Bansal et al., 2023c, a; Motwani and Brun, 2023) 6
@@
String Replacement (Nashid et al., 2023; Xu et al., 2022; Zhou et al., 2019; Wang et al., 2021a; Bansal et al., 2023a; Zügner et al., 2021; Lin et al., 2021) 7
Number Replacement (Nashid et al., 2023; Xu et al., 2022; Wang et al., 2021a; Zügner et al., 2021; Lin et al., 2021) 5
@@
Other Operator (Chen et al., 2021a; McBurney and McMillan, 2016; Kim et al., 2019; Yu et al., 2023; Zou et al., 2020; Nie et al., 2021; Wang et al., 2021b; Vu et al., 2023; Phan et al., 2024; Bi et al., 2024b; Lyu et al., 2021; Nashid et al., 2024; Liang et al., 2024; Liao et al., 2024; Xie et al., 2024; Cheng et al., 2024; Li et al., 2024d; Wang et al., 2024f; Mandli et al., 2024; Wen et al., 2024b; Xu et al., 2021; Yang et al., 2024b; Wang et al., 2024d; Liu et al., 2024b; Lu et al., 2022; Bibaev et al., 2022; Wang et al., 2023; Li et al., 2024b; Jiang et al., 2024; Nafi et al., 2019, 2018; Zhao et al., 2024; Abdu et al., 2024; Zhou et al., 2024; Zhang et al., 2024a) 35

### v2

@@ 3.3. Inclusion and Exclusion Criteria
After identifying studies that align with our research objectives, it is necessary to filter out unqualified studies. To achieve this, we apply inclusion and exclusion criteria to assess the quality of candidate studies. Each retained research must satisfy all inclusion and exclusion criteria. This evaluation ensures that each retained research primarily focuses on using context to address one type of code intelligence task. The inclusion and exclusion criteria are as follows.
@@ 3.4. Data Extraction and Collection
After filtering out irrelevant and duplicate papers, as shown in Figure 1, we extract relevant data and analyze it to address our four research questions. For RQ1, we record the year, type, publication venue, primary contributions, and task focus of each study. These details are vital for mapping the evolution of the research landscape, revealing historical trends, and understanding how various publication venues have influenced the development of context-aware code intelligence. For RQ2, we examine the types of context used and analyze trends in their application across seven code intelligence tasks. Collecting these data points is instrumental in identifying the diversity of context usage and in assessing its potential application scenarios. For RQ3, we review context preprocessing methods and context-aware models introduced in papers that present new techniques. This data collection sheds light on the methodological innovations and technical challenges involved in effectively leveraging context, which is crucial for advancing model development. For RQ4, we collate evaluation metric(s), dataset(s) created by the paper (excluding datasets created from external sources), and open-source status (availability of code, model weights, and datasets created by the paper).
cross-verification process for our collected data. The set of 146 papers was divided between two researchers (the paper’s authors), who independently read their assigned papers and populated the extraction schema. Upon completion, the two researchers swapped their entire completed sets. Each researcher then performed a full verification of the other’s work, checking every extracted data point against the original source paper. Any discrepancies were flagged and resolved through discussion and joint re-examination of the source paper by the two researchers. A third, senior author was designated as the arbiter to resolve any remaining disagreements, ensuring consistency across the entire dataset.
@@ 4. RQ1: What Are the Trends and Distributions of the Publications Studied?
To understand the development trends of papers that utilize context in code intelligence, we analyze the trends and distributions of the reviewed publications. Our analysis focuses on three key aspects: publication dates, venues, and main contributions of the primary studies.
@@ 4.1. Trends of Publication Year
Refer to caption Short description Number of Papers Published Per Year
@@
Although the concept of “context” had been explored in natural language processing tasks before 2007 (Porzel and Gurevych, 2003), to the best of our knowledge, we did not find any papers prior to 2007 that used context to address code intelligence tasks in our investigation. Therefore, our analysis focuses on relevant papers published from 2007 to 2024. Figure  2 illustrates the publication trend of works employing context over the past eighteen years. To provide a clearer view of the trend in context utilization within code intelligence, we divide the research of these papers presented in these papers into three distinct phases.
@@
The second phase is from 2019 to 2022, during which more relevant work emerges. The average number of published papers per year rises to 9.5. This increase is related to earlier work that laid the foundation for context utilization and developing deep learning models, such as the transformer architecture (Vaswani et al., 2017). At this stage, all seven types of code intelligence tasks begin to investigate how to fully utilize contextual information.
The third phase is from 2023 to 2024. In this phase, ChatGPT (Wu et al., 2023) initiates a wave of large language models, whose outstanding performance drives code intelligence tasks to further explore the use of context in large language models. During this phase, the number of papers increases substantially, with 31 papers published in 2023 and 49 in 2024 (as of August). Except for the clone detection task, all tasks show substantial growth in the number of papers. We suspect that clone detection may be less dependent on large language models.
Refer to caption Short description (Left) Number of Publications Per Type. (Right) Number of Publications Per Venue.
@@ 4.3. Types of Main Contributions of the Studied Papers
Refer to caption Short description (Left) Number of Papers Per Type of Main Contribution. (Right) Number of Papers Per Tasks.
@@
We simultaneously analyze the proportions of each type of code intelligence task, as shown in Figure 4. Top-3 code intelligence tasks are defect detection (19.9%), code completion (19.2%), and code summarization (16.4%), while code generation occupies 15.8% and program repair contains 13.7%. Commit message generation and clone detection take the minor parts, 8.2%, and 6.8%, respectively.
@@ 5. RQ2: How Are Contexts Categorized in Code Intelligence Tasks?
To categorize context in code intelligence, we divide it into two types based on how it is obtained: direct and indirect. Direct context refers to information that can be accessed without further processing, such as source code, API documentation, and commit messages. Indirect context, however, requires additional processing or transformation, examples of which include abstract syntax trees (ASTs), control flow graphs (CFGs), and compiler-generated information. This section will explore the different context categories and examine trends in their use in code intelligence tasks.
@@ 5.1. Types of Context
As shown in Figure  5, we summarize all types of contexts into a tree diagram. The number after the context name indicates how many new technique papers use this context, while benchmark, case study, and empirical papers are excluded. The combined count of direct context utilization (97 papers) and indirect context utilization (51 papers) exceeds the total number of papers (122 papers) because 26 papers utilize both direct and indirect contexts. Overall, direct contexts are more prevalent than indirect contexts. However, researchers discovered that transforming direct context into various indirect contexts in specific scenarios can improve model performance and efficiency. For example, Zhang et al. (Zhang et al., 2023b) propose a graph neural network based on control flow graphs to detect condition-related bugs, which significantly outperforms state-of-the-art methods. Li et al. (Li, 2020) embed abstract syntax trees, data flow graphs, and program dependency graphs to capture local and global context for bug detection. These representations outperform the other representations by up to 206% in accuracy.
Refer to caption Short description Types of Context Used in the Seven Studied Code Intelligence Tasks
@@
Refer to caption Short description Number of Context Utilized in Each Code Intelligence Task
Figure 6. Number of Context Utilized in Each Code Intelligence Task
@@ 5.1.1. Direct Context
Eighty papers adapt code-based contexts, which are the most numerous of all forms of contexts, including sixty-eight papers for source code and seventeen papers for code diffs. Source code, the most popular form of context, constructs the basis for program execution. Mining and extracting source code context mitigates hallucination and provides more accurate answers by offering models with specific domain knowledge in repositories (Nashid et al., 2024; Zhang et al., 2023a). Code diff context presents changes in source code over time or across versions, providing more fine-grained information than source code context in tasks such as defect detection (Youm et al., 2017) and commit message generation (Li et al., 2024d).
@@ 5.1.2. Indirect Context
Indirect context can be categorized into three types: tree-based, graph-based, and sequence-based contexts. Tree-based context includes AST (24 papers), and constructing ASTs showcases the structure of code. The graph-based context includes CFG (9 papers), DFG (6 papers), PDG (6 papers), CPG (5 papers), and UML (2 papers), providing important information about source code, such as data flow, control flow, logic, and structure. Sequence-based context includes compilation information (9 papers) and IDE information (2 papers). Compilation information offers insight into the dynamics of code execution, and IDE information provides information such as programmers’ usage patterns.
@@ 5.2. Trends of Context
Besides, we find that source code is the most popular form of context, utilized by all seven code intelligence tasks, followed by AST (utilized by six tasks) and CFG (utilized by five tasks). In all 13 kinds of context, defect detection employs the widest variety of context (9 types), while program repair and clone detection both select eight kinds of context, respectively. The variation in the amount of context utilized across different code intelligence tasks is due to how well the context information aligns with the specific requirements of each task.
However, some kinds of context are less common. Commit message and IDE information are utilized twice by commit message generation (Huang et al., 2017; Eliseeva et al., 2023) and code completion (Li et al., 2024e; Bibaev et al., 2022), respectively. UML is selected by code generation (E. V. and Samuel, 2019) and code summarization (Wang et al., 2021b). The limited application of these contexts is largely attributable to the scarcity of publicly available, standardized datasets and the inherent difficulty in translating their information for specific tasks. Besides, the reasons, such as data privacy for IDE logs and the high abstraction level of UML diagrams, also limit the usage of this context.
@@ 6. RQ3: What Methods Are Used to Preprocess and Model Context?
Refer to caption Short description The Pipeline of Context Utilization
Figure 7. The Pipeline of Context Utilization
While context provides valuable information, it must be carefully processed and modeled to be effectively used in code intelligence tasks. Preprocessing refers to the techniques applied to the raw context data to structure and transform it into a usable input representation. Modeling involves the architectural workflow and methods used to integrate this already-preprocessed context representation into models to inform and improve task performance. This section examines how researchers process context, focusing on both preprocessing methods and modeling strategies used in various code intelligence tasks. Figure 7 illustrates the end-to-end pipeline of context utilization across code intelligence tasks, including context mining, context preprocessing, and context modeling.
@@ 6.1. Context Preprocess
To intuitively understand the effect of context preprocessing in each paper, we categorize the preprocessing operators into four types: splitting, irrelevance removal, unifying, and other operators. The details of the preprocessing steps are shown in Table  4.
@@ 6.1.1. Splitting
Byte pair encoding (BPE) (Gage, 1994) is a widely used splitting method in natural language processing. BPE reduces the number of tokens by iteratively converting frequently occurring data into new units, alleviating the training burden while maintaining data accuracy. For example, Liu et al. (Liu et al., 2023a) employ BPE method to segment both natural language and source code data, facilitating the generation of feature embeddings. This technique can be applied in code intelligence tasks to process code-relevant context, such as API documentation (Zan et al., 2023, 2022; Zhou et al., 2023), compiler information (Agrawal et al., 2024), code comments (Liu et al., 2023a), source code contexts (Ding et al., 2023b; Shrivastava et al., 2023b; Zhang et al., 2023a).
Programming languages are highly structured and adhere to stricter syntax rules than natural language descriptions. To facilitate program comprehension, programmers often adopt CamelCase or SnakeCase as naming conventions for coding identifiers, such as variables, functions, and classes. These conventions offer practical guidelines for splitting identifiers: CamelCase identifiers can be divided by detecting transitions between lowercase and uppercase letters, while SnakeCase identifiers are split at underscores (’_’). Both techniques are widely used and effective in preprocessing source code context (Lutellier et al., 2020; Wang et al., 2021b; Bansal et al., 2023c). For example, Wang et al. (Wang et al., 2021b) employ the CamelCase rule to split class names in UML, while Bansal (Bansal et al., 2023c) utilizes the CamelCase rule to split cross-file source code tokens.
Additionally, segmentation rules based on non-alphabetic (Washio and Miyao, 2022; Xu et al., 2022; Gao et al., 2023) or non-numeric symbols (Washio and Miyao, 2022; Shrivastava et al., 2023a) can serve as supplementary approaches for segmenting both code and natural language content. These methods are widely used for their simplicity, as they do not rely on complex lexicons or language models and serve as an initial approach to token segmentation.
Another approach is to use a tokenizer from an established library or provided by language models. Examples of established libraries include NLTK   [^3] https://github.com/nltk/nltk (Lin et al., 2021; Bangcharoensap et al., 2012), Javalang   [^4] https://github.com/c2nes/javalang (Zhang et al., 2023b) and Pygments   [^5] https://github.com/pygments/pygments (Zügner et al., 2021). Examples of tokenizers provided by language models include those from StarCoder (Phan et al., 2024; Cheng et al., 2024; Liu et al., 2024b), GPT-4 (Xie et al., 2024), and RoBERTa (Wang et al., 2024b). A tokenizer segments continuous text into discrete tokens, assigns each token a unique index, and maps the index to token embeddings for further processing.
@@
Splitting BPE (Agashe et al., 2019; Agrawal et al., 2024; Zan et al., 2023, 2022; Tang et al., 2023; Zhou et al., 2023; Ding et al., 2023b; Shrivastava et al., 2023b; Zhang et al., 2023a; Liu et al., 2023a; Ding et al., 2023a; Qiu et al., 2024) 12
Camel_Case (Lutellier et al., 2020; Lin et al., 2022; Saha et al., 2013; Yu et al., 2020; Shahbazi and Fard, 2023; Shahbazi et al., 2021; Guo et al., 2023b; Shrivastava et al., 2023b; Wang et al., 2021b; Bansal et al., 2023c, a; Zügner et al., 2021; Huang et al., 2017; Lin et al., 2021; Bangcharoensap et al., 2012; Youm et al., 2015; Swe and Oo, 2019; Zhou et al., 2012; Jiang et al., 2018; Wang and Lo, 2014; Gao et al., 2023; Wong et al., 2015; Motwani and Brun, 2023; Youm et al., 2017; Zhang et al., 2021) 25
Snake_Case (Lin et al., 2022; Yu et al., 2020; Shahbazi and Fard, 2023; Shahbazi et al., 2021; Guo et al., 2023b; Shrivastava et al., 2023b; Wang et al., 2021b; Bansal et al., 2023c, a; Zügner et al., 2021; Zhang et al., 2021) 11
@@
Base on Tokenizer (Zügner et al., 2021; Lin et al., 2021; Zhang et al., 2023b; Bangcharoensap et al., 2012; Phan et al., 2024; Xie et al., 2024; Wang et al., 2024a; Cheng et al., 2024; Wang et al., 2024b; Liu et al., 2024b; Deng et al., 2024) 11
@@
Punctuations Filtering (Shahbazi and Fard, 2023; Shahbazi et al., 2021; Wang et al., 2021b; Bansal et al., 2023a; Koyuncu et al., 2019; Matmti and Fard, 2023; Wang and Lo, 2014; Motwani and Brun, 2023; Youm et al., 2017; Zhang et al., 2021) 10
Comment Removal (Zügner et al., 2021; Guo et al., 2024b; Imani et al., 2024) 1
@@
Symbol Removal (Guo et al., 2024b) 1
Unifying Lowercase (Shahbazi and Fard, 2023; Shahbazi et al., 2021; Wang et al., 2021b; Bansal et al., 2023c, a; Motwani and Brun, 2023) 6
@@
String Replacement (Nashid et al., 2023; Xu et al., 2022; Zhou et al., 2019; Wang et al., 2021b; Bansal et al., 2023a; Zügner et al., 2021; Lin et al., 2021) 7
Number Replacement (Nashid et al., 2023; Xu et al., 2022; Wang et al., 2021b; Zügner et al., 2021; Lin et al., 2021) 5
@@
Other Operator (Chen et al., 2021b; McBurney and McMillan, 2016; Kim et al., 2019; Yu et al., 2023; Zou et al., 2020; Nie et al., 2021; Wang et al., 2021a; Vu et al., 2023; Phan et al., 2024; Bi et al., 2024; Lyu et al., 2021; Nashid et al., 2024; Liang et al., 2024; Liao et al., 2024; Xie et al., 2024; Cheng et al., 2024; Li et al., 2024e; Wang et al., 2024b; Mandli et al., 2024; Wen et al., 2024; Xu et al., 2021; Yang et al., 2024; Wang et al., 2024g; Liu et al., 2024b; Lu et al., 2022; Bibaev et al., 2022; Wang et al., 2023; Li et al., 2024a; Jiang et al., 2024; Nafi et al., 2019, 2018; Zhao et al., 2024; Abdu et al., 2024; Zhou et al., 2024; Zhang et al., 2024d) 35